OpenStudy (anonymous):

Solve the exponential equation. Express the solution in terms of natural logarithms. Then use a calc. to obtain a decimal approximation for the solution. 4^(x+)= 301

OpenStudy (anonymous):

x plus what?

OpenStudy (anonymous):

oh sorry it's x+3 :)

OpenStudy (anonymous):

@AnonymousBeast do you think you can still maybe help me out? It would be greatly appreciated :)

OpenStudy (anonymous):

@D3xt3R Could you please help me?

OpenStudy (anonymous):

Yes, I'm here.

OpenStudy (anonymous):

Rewrite in logarithmic form. Use this method: \[a^b=y \rightarrow \log_{a}y=b \]

OpenStudy (anonymous):

\[4^{x+3}=301\]we can write\[4^{x+3}=2^{2*(x+3)}=2^{2x+6}\]now writing with the new form\[2^{2x+6}=301\]apply natural logarithms on both sides\[ln(2^{2x+6})=ln(301)\]\[(2x+6)*ln(2)=ln(301)\]\[2x+6=\frac{ln(301)}{ln(2)}\]\[2x+6=log_{_2}(301)\]\[2x=log_{_2}(301)-6\]\[\boxed{\boxed{x=\frac{log_{_2}(301)-6}{2}}}\]

OpenStudy (anonymous):

Or that works too :P

OpenStudy (anonymous):

This stuff is so confusing!! Do you think you can explain it a little more maybe??

OpenStudy (anonymous):

I can show you another way perhaps?

OpenStudy (anonymous):

I just apply a logarithm property, and it is.\[log_c(a^b)\Leftrightarrow b*log_c(a)\]

OpenStudy (anonymous):

So what exactly do I need to do? I am soo sorry guys! :/

OpenStudy (anonymous):

After converting to logarithmic form, use the change of base rule: \[\log_{4} 301=\frac{ \log 301 }{ \log 4 }\] Use a calculator. It is of base 10, so just about all scientific calculators can solve. You get about 4.1168098383798510231630554788499 Now we have: 4.1168098383798510231630554788499=x+3 Simplify: x= approximately 1.1168098383798510231630554788499

OpenStudy (anonymous):

Oh, my bad. While x is correct, I did not see to express it in NATURAL logarithms. Sorry :/

OpenStudy (anonymous):

This is Algebra II, right?

OpenStudy (anonymous):

It's okay :) How do you express it in natural logarithms?

OpenStudy (anonymous):

D3xter expressed it in terms of ln. His result should be your answer I believe.

OpenStudy (anonymous):

\[natural~logarithms~=~ln(anything)~=~log_{_e}(anything)\]

OpenStudy (anonymous):

So, I am so confused. It's not your answer @AnonymousBeast

OpenStudy (anonymous):

What are your choices?

OpenStudy (anonymous):

We don't have choices :/

OpenStudy (anonymous):

If you substitute about 1.1168098383798510231630554788499 for x into the original problem, it'll get to about 100.99... which is about as close as you'll get. Like I said, I can't express this in natural logarithms. @D3xt3R may be able to help you.

OpenStudy (anonymous):

Guy's I'm from Brazil, can you tell me if you can use calculator??

OpenStudy (anonymous):

I think you can

OpenStudy (anonymous):

Can you please express what AnonymousBeast said into natural logarithms??

OpenStudy (anonymous):

You would have to in order to do logarithms, really.

OpenStudy (anonymous):

Do I express 100.99 in natural log?

OpenStudy (anonymous):

@CookieMonsta22 first step, choose witch log do you wanna you, second step apply this log in both sides, third step use the properties of logarithms, then you will find the value of X, like i did, now using a calculator you have to write this sentence to find the value\[x=\frac{ln(301)}{ln(2)}-6\]you'll find\[\boxed{\boxed{x\approx2.2336}}\]

OpenStudy (anonymous):

So is x=2.2336 the final answer?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Oh my gosh thank you so so so much!!! Both of you are amazing! I cannot thank you enough :) <3

OpenStudy (anonymous):

How do I award you guys medals??

OpenStudy (anonymous):

@AnonymousBeast and @D3xt3R how do I award you medals?

OpenStudy (anonymous):

Do you guys know what the solution in terms of natural logarithms is? It asks for the solution set??

OpenStudy (anonymous):

The solution of a natural logarithms is a rational number or an irrational number

OpenStudy (anonymous):

@CookieMonsta22 I read you question again, and the right answer is \[x=\frac{ln(301)-6~*~ln(2)}{ln(2)}\]you have to do the minimum common multiple (MCM), and don't have to disappear with the natural logarthms.

OpenStudy (anonymous):

Is that the solution set @D3xt3R

OpenStudy (anonymous):

Yes, you don't have to replace the real value of ln(301) and ln(2), you just have to answer like I did.

OpenStudy (anonymous):

Okay so that's the final answer for the solution set?

OpenStudy (anonymous):

@D3xt3R so that's the final answer for the solution set?

OpenStudy (mathstudent55):